OpenStudy (anonymous):

Xsqrt3=12 What is X?

OpenStudy (anonymous):

How would I work it out, or find the answer?

OpenStudy (anonymous):

x=12/sqrt3

OpenStudy (anonymous):

6.9?

OpenStudy (anonymous):

Basically, you solve for x x√3 = 12 x = 12/√3 x = 12√3/3 x = 4√3

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

x = 4sqrt3 for a more purrtty answer

OpenStudy (anonymous):

Where did y'all get the 4? @Calcmathlete

OpenStudy (anonymous):

It's something called rationalizing the denominator. You multiply both the numerator and denominator by √3 to get the √3 out of the denominator. Then, 12 and 3 both have 3 as a common factor, so you simplify to get 4. 12/3 = 4. 3/3 = 1 which is invisible.